# Scaling Laws Meet Model Architecture: Toward Inference-Efficient LLMs

### v1

@@ 6 Related Work
Scaling Laws for Language Models.
Scaling laws are powerful tools to predict the performance of large language models. Existing scaling laws Hoffmann et al. (2022); Muennighoff et al. (2023); Sardana et al. (2023); Kumar et al. (2024); Gadre et al. (2024); Ruan et al. (2024) characterize how model performance varies with model size, dataset size, data quality, and compute budget. With the rise of Mixture-of-Experts (MoE) Shazeer et al. (2017); Guo et al. (2025), a powerful architecture for large language models, recent studies Krajewski et al. (2024); Abnar et al. (2025) extend scaling laws to account for the number of experts, expert granularity, active parameters, and sparsity.
@@
Efforts to improve the inference efficiency of large language models generally fall into two categories: one line of work investigates the trade-offs across different model configurations Alabdulmohsin et al. (2023); Bian et al. (2025), while the other focuses on designing more efficient model architectures Xiao et al. (2023); Gu & Dao (2023); Gao et al. (2024b); Jiang et al. (2024); Liu et al. (2024b); Dao & Gu (2024); Xiao et al. (2024); Yuan et al. (2025); Chandrasegaran et al. (2025).
7 Limitations and Future Work
While our team has made notable progress, several open challenges remain that offer promising directions for future research. First, due to limitations in resources and time, our evaluation does not extend to 7B models. Second, our analysis is restricted to dense models, and it remains unclear whether the results extend to Mixture of Experts (MoE) architectures Shazeer et al. (2017). While we report inference efficiency measurements for MoE models under varying architectural choices in Appendix J, we have not yet established scaling laws for MoE architectures. Third, we adopt the experimental setup from Chen et al. (2025), and it is uncertain whether different model architectures warrant different hyperparameter configurations. Finally, our analysis is limited to pre-training, and it remains unclear how the results would change under post-training.
8 Conclusion
This work explores the trade-off between model accuracy and inference cost under a fixed training budget. We begin by demonstrating how architectural choices influence both inference throughput and model accuracy. Building on this, we extend Chinchilla scaling laws to incorporate architectural factors and propose a framework for optimal model architecture search. Using the fitted scaling laws and our framework, we trained models up to 3B parameters, achieving up to 42% higher inference throughput and 2.1% accuracy gains across nine downstream tasks.
Reproducibility Statement
All experiments in this work were conducted using publicly available frameworks. Section 4 provides details of our training, inference, and evaluation setups. In particular, we used Megatron-LM (Shoeybi et al., 2019) for model training, vLLM (Kwon et al., 2023) for efficient inference, and lm-eval-harness (Gao et al., 2024a) for standardized evaluations. To facilitate reproducibility, we will release configuration files and scripts.
Appendix A LLM Usage
Appendix B Open-Weighted Model Architectures
@@
Table 3: Open-Weighted Model Architectures: We list the architectural configurations of all models used in this paper. $n_{\text{layers}}$ is the number of layers, $d_{\text{model}}$ is the hidden size, $n_{\text{heads}}$ is the number of attention heads, and $f_{\text{size}}$ is the intermediate size.
@@
Appendix C Model Architectures
@@
Table 4: Model Architectures: We list the architectural configurations of all models trained in this paper. $N_{\text{non-embed}}$ is the total number of non-embedding parameters, $n_{\text{layers}}$ is the number of layers, $d_{\text{model}}$ is the hidden size, $n_{\text{heads}}$ is the number of attention heads, $f_{\text{size}}$ is the intermediate size, and $r_{\text{mlp}/\text{attn}}$ is the MLP-to-attention ratio.
@@
Appendix D Hyper-parameters
@@
Table 5: Hyper-parameters: We show the hyper-parameters used for training in this paper.
@@
Appendix E Additional Inference Evaluation Results
@@
In this section, we present additional inference efficiency results on NVIDIA A100 GPUs. Figure 10 presents that, when parameter count, MLP-to-Attention ratio, and hidden size are fixed, increasing GQA leads to higher inference throughput, consistent with the findings of Ainslie et al. (2023). We alter model configurations of LLaMA-3.2-1B, 3B, and LLaMA-3.1-8B in Figure 10.
@@
Figure 8: Hidden size on Inference Throughput: (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. Across varying batch sizes and model scales, larger hidden sizes yield higher inference throughput under a fixed parameter budget. The legend indicates the hidden size of the models, where $d=d_{\text{model}}$.
@@
Figure 9: MLP-to-Attention ratio on Inference Throughput: (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. Across varying batch sizes and model scales, a larger MLP-to-Attention ratio increases inference throughput under a fixed parameter budget. The legend indicates the MLP-to-Attention ratio of the models, where $r=r_{\text{mlp}/\text{attn}}$.
@@
Figure 10: GQA on Inference Throughput: (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. This figure shows the impact of GQA on inference throughput. With the total parameter count fixed, hidden size is set to 2048 (1B), 3072 (3B), and 4096 (8B), and the MLP-to-Attention ratio is 4.0, 2.67, and 4.2, respectively. Across varying batch sizes, models with larger GQA achieve higher throughput. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
@@
Furthermore, we derive architectural variants by altering the configurations of Qwen3-0.6B, 1.7B, and 4B to investigate the impact of model architectural factors on inference efficiency. The results are shown in Figure 11-13.
@@
Figure 11: Hidden size on Inference Throughput (Qwen3): (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, larger hidden sizes yield higher inference throughput under a fixed parameter budget. The legend indicates the hidden size of the models, where $d=d_{\text{model}}$. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
@@
Figure 12: MLP-to-Attention ratio on Inference Throughput (Qwen3): (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, a larger MLP-to-Attention ratio increases inference throughput under a fixed parameter budget. The legend indicates the MLP-to-Attention ratio of the models, where $r=r_{\text{mlp}/\text{attn}}$. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
@@
Figure 13: GQA on Inference Throughput (Qwen3): (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. This figure shows the impact of GQA on inference throughput. With the total parameter count fixed, hidden size is set to 1024 (0.6B), 2048 (1.7B), and 2560 (4B), and the MLP-to-Attention ratio is 1.5, 3.0, and 2.85, respectively. Across varying batch sizes, models with larger GQA achieve higher throughput. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
Appendix F Additional Results: Loss vs. Model Architecture
In this section, we analyze the relationship between training loss and GQA while fixing the number of parameters, hidden size, and MLP-to-Attention ratio. As shown in Figure 14, unlike hidden size and MLP-to-Attention ratio, the relationship between loss and GQA is highly fluctuating.

### v2

@@ Appendix B Additional Related Work
Efforts to improve the inference efficiency of large language models generally fall into two categories: one line of work investigates the trade-offs across different model configurations Alabdulmohsin et al. (2023); Bian et al. (2025); Bian (2026), while the other focuses on designing more efficient model architectures Xiao et al. (2023); Gu and Dao (2023); Gao et al. (2024b); Jiang et al. (2024); Liu et al. (2024b); Dao and Gu (2024); Xiao et al. (2024); Yuan et al. (2025); Chandrasegaran et al. (2025).
Appendix C Open-Weighted Model Architectures
@@
Table 3: Open-Weighted Model Architectures. We list the architectural configurations of all models used in this paper. $n_{\text{layers}}$ is the number of layers, $d_{\text{model}}$ is the hidden size, $n_{\text{heads}}$ is the number of attention heads, and $f_{\text{size}}$ is the intermediate size.
@@
Appendix D Model Architectures
@@
Table 4: Model Architectures. We list the architectural configurations of all models trained in this paper. $N_{\text{non-embed}}$ is the total number of non-embedding parameters, $n_{\text{layers}}$ is the number of layers, $d_{\text{model}}$ is the hidden size, $n_{\text{heads}}$ is the number of attention heads, $f_{\text{size}}$ is the intermediate size, and $r_{\text{mlp}/\text{attn}}$ is the MLP-to-attention ratio.
@@
Appendix E Hyper-parameters
@@
Table 5: Hyper-parameters. We show the hyper-parameters used for training in this paper.
@@
Appendix F Additional Inference Evaluation Results over A100 GPUs
@@
In this section, we present additional inference efficiency results on NVIDIA A100 GPUs. Figure 11 presents that, when parameter count, MLP-to-Attention ratio, and hidden size are fixed, increasing GQA leads to higher inference throughput, consistent with the findings of Ainslie et al. (2023). We alter model configurations of LLaMA-3.2-1B, 3B, and LLaMA-3.1-8B in Figure 11. Moreover, we use the SGLang framework Zheng et al. (2023) to benchmark the inference throughput of LLaMA-3.2-1B, LLaMA-3.2-3B, Surefire-1B, and Surefire-3B on a single A100 GPU in Figure 15.
@@
Figure 9: Hidden size on Inference Throughput. (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. Across varying batch sizes and model scales, larger hidden sizes yield higher inference throughput under a fixed parameter budget. The legend indicates the hidden size of the models, where $d=d_{\text{model}}$.
@@
Figure 10: MLP-to-Attention ratio on Inference Throughput. (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. Across varying batch sizes and model scales, a larger MLP-to-Attention ratio increases inference throughput under a fixed parameter budget. The legend indicates the MLP-to-Attention ratio of the models, where $r=r_{\text{mlp}/\text{attn}}$.
@@
Figure 11: GQA on Inference Throughput. (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. This figure shows the impact of GQA on inference throughput. With the total parameter count fixed, hidden size is set to 2048 (1B), 3072 (3B), and 4096 (8B), and the MLP-to-Attention ratio is 4.0, 2.67, and 4.2, respectively. Across varying batch sizes, models with larger GQA achieve higher throughput. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
@@
Furthermore, we derive architectural variants by altering the configurations of Qwen3-0.6B, 1.7B, and 4B to investigate the impact of model architectural factors on inference efficiency. The results are shown in Figure 12-14.
@@
Figure 12: Hidden size on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, larger hidden sizes yield higher inference throughput under a fixed parameter budget. The legend indicates the hidden size of the models, where $d=d_{\text{model}}$. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
@@
Figure 13: MLP-to-Attention ratio on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, a larger MLP-to-Attention ratio increases inference throughput under a fixed parameter budget. The legend indicates the MLP-to-Attention ratio of the models, where $r=r_{\text{mlp}/\text{attn}}$. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
@@
Figure 14: GQA on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. This figure shows the impact of GQA on inference throughput. With the total parameter count fixed, hidden size is set to 1024 (0.6B), 2048 (1.7B), and 2560 (4B), and the MLP-to-Attention ratio is 1.5, 3.0, and 2.85, respectively. Across varying batch sizes, models with larger GQA achieve higher throughput. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
Figure 15: Results for 1B and 3B models. (left) Inference throughput comparison between LLaMA-3.2-1B and Surefire-1B, showing that Surefire-1B consistently achieves higher efficiency across batch sizes. (right) Inference throughput comparison between LLaMA-3.2-3B and Surefire-3B, demonstrating that Surefire-3B consistently delivers higher efficiency across all batch sizes. The results are collected using the SGLang framework Zheng et al. (2023) on a single A100 GPU with 4096 input and 1024 output tokens.
Appendix G Additional Inference Evaluation Results over H200 GPUs
In this section, we present additional inference efficiency results on NVIDIA H200 GPUs. We derive architectural variants by altering the configurations of Qwen3-0.6B, 1.7B, and 4B to investigate the impact of model architectural factors on inference efficiency. The results are shown in Figure 16-18. We also compare the inference throughput of Surefire-1B, Surefire-3B, with LLaMA-3.2-1B and LLaMA-3.2-3B over H200 GPUs in Figure 19.
Figure 16: Hidden size on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, larger hidden sizes yield higher inference throughput under a fixed parameter budget. The legend indicates the hidden size of the models, where $d=d_{\text{model}}$. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA H200 GPU with 4096 input and 1024 output tokens.
Figure 17: MLP-to-Attention ratio on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, a larger MLP-to-Attention ratio increases inference throughput under a fixed parameter budget. The legend indicates the MLP-to-Attention ratio of the models, where $r=r_{\text{mlp}/\text{attn}}$. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA H200 GPU with 4096 input and 1024 output tokens.
Figure 18: GQA on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. This figure shows the impact of GQA on inference throughput. With the total parameter count fixed, hidden size is set to 1024 (0.6B), 2048 (1.7B), and 2560 (4B), and the MLP-to-Attention ratio is 1.5, 3.0, and 2.85, respectively. Across varying batch sizes, models with larger GQA achieve higher throughput. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA H200 GPU with 4096 input and 1024 output tokens.
Figure 19: Results for 1B and 3B models. (left) Inference throughput comparison between LLaMA-3.2-1B and Surefire-1B, showing that Surefire-1B consistently achieves higher efficiency across batch sizes. (right) Inference throughput comparison between LLaMA-3.2-3B and Surefire-3B, demonstrating that Surefire-3B consistently delivers higher efficiency across all batch sizes. The results are collected using the SGLang framework Zheng et al. (2023) on a single NVIDIA H200 GPU with 4096 input and 1024 output tokens.
Furthermore, we use the SGLang framework Zheng et al. (2023) to measure the inference throughput of large language models. We construct architectural variants by modifying the configurations of Qwen3-0.6B, 1.7B, and 4B to study how different architectural factors influence inference efficiency. The results are presented in Figure 20-22. The inference throughput of Surefire-1B and Surefire-3B compared with LLaMA-3.2-1B and LLaMA-3.2-3B on H200 GPUs is shown in Figure 23.
Figure 20: Hidden size on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, larger hidden sizes yield higher inference throughput under a fixed parameter budget. The legend indicates the hidden size of the models, where $d=d_{\text{model}}$. All evaluations are performed using the SGLang framework Zheng et al. (2023) on a single NVIDIA H200 GPU with 4096 input and 1024 output tokens.
Figure 21: MLP-to-Attention ratio on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, a larger MLP-to-Attention ratio increases inference throughput under a fixed parameter budget. The legend indicates the MLP-to-Attention ratio of the models, where $r=r_{\text{mlp}/\text{attn}}$. All evaluations are performed using the SGLang framework Zheng et al. (2023) on a single NVIDIA H200 GPU with 4096 input and 1024 output tokens.
Figure 22: GQA on Inference Throughput (Qwen3). (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. This figure shows the impact of GQA on inference throughput. With the total parameter count fixed, hidden size is set to 1024 (0.6B), 2048 (1.7B), and 2560 (4B), and the MLP-to-Attention ratio is 1.5, 3.0, and 2.85, respectively. Across varying batch sizes, models with larger GQA achieve higher throughput. All evaluations are performed using the SGLang framework Zheng et al. (2023) on a single NVIDIA H200 GPU with 4096 input and 1024 output tokens.
Figure 23: Results for 1B and 3B models. (left) Inference throughput comparison between LLaMA-3.2-1B and Surefire-1B, showing that Surefire-1B consistently achieves higher efficiency across batch sizes. (right) Inference throughput comparison between LLaMA-3.2-3B and Surefire-3B, demonstrating that Surefire-3B consistently delivers higher efficiency across all batch sizes. The results are collected using the SGLang framework Zheng et al. (2023) on a single NVIDIA H200 GPU with 4096 input and 1024 output tokens.
Appendix H Detailed Throughput Statistics
In this section, we present the detailed inference throughput results for the 1B and 3B models in Table 6.
Table 6: Summary of Results for 1B and 3B Models. We summarize the inference throughput (tokens/s) of LLaMA-3.2-1B, Surefire-1B, LLaMA-3.2-3B, and Surefire-3B across vLLM and SGLang on A100 and H200 GPUs using 4096 input tokens and 1024 output tokens.
Hardware Framework Model Batch Size
16 32 64 128
A100 vLLM LLaMA-3.2-1B 1931.87 2602.72 3409.85 3825.91
Surefire-1B 2116.49 3290.23 4028.69 4800.05
LLaMA-3.2-3B 904.83 1121.39 1136.61 1222.03
Surefire-3B 1005.44 1356.07 1613.32 1476.22
A100 SGLang LLaMA-3.2-1B 2748.84 3643.27 4703.92 5353.29
Surefire-1B 3239.55 4737.63 5832.01 6962.24
LLaMA-3.2-3B 1173.51 1452.97 1668.67 1762.18
Surefire-3B 1318.23 1726.20 2081.44 2251.74
H200 vLLM LLaMA-3.2-1B 4311.97 6221.14 8131.65 9306.36
Surefire-1B 4532.85 6992.71 9493.46 11282.56
LLaMA-3.2-3B 2269.53 3119.94 3872.14 4311.43
Surefire-3B 2309.48 3271.63 4242.33 4841.53
H200 SGLang LLaMA-3.2-1B 4812.67 6939.88 8038.34 8608.57
Surefire-1B 5900.52 8798.68 11214.40 12645.55
LLaMA-3.2-3B 2593.04 3370.42 3868.42 4183.09
Surefire-3B 2542.21 3488.79 4446.66 4877.16
We further compare design choices across existing open-source models at the 1B and 3B scales in Table 7 and Table 8. For the LLaMA-3.2-1B, Panda-1B, and Surefire-1B models we pretrained, we report inference throughput (tokens/s), byte-level WikiText perplexity, and full architectural configurations in the accompanying tables. All throughput measurements are performed with vLLM on H200 GPUs using batch size 128. For the 1B scale, we include LLaMA-3.2-1B-HF and OLMo-2-1B-HF. Because OLMo supports only a 4k context window and cannot run our standard 4k/1k setup (4096 input tokens and 1024 output tokens), we additionally report results under a 2k/1k setup (2048 input tokens and 1024 output tokens). For the 3B scale, we add LLaMA-3.2-3B-HF and Qwen2.5-3B-HF, all evaluated under the 4k/1k configuration.
Table 7: Comparison against open-source models at the 1B scale. We compare our pretrained LLaMA-3.2-1B, Panda-1B, and Surefire-1B models with LLaMA-3.2-1B-HF and OLMo-2-1B-HF in terms of inference throughput (on H200 GPUs using vLLM) and byte-level WikiText perplexity.
Model LLaMA-3.2-1B Panda-1B Surefire-1B LLaMA-3.2-1B-HF OLMo-2-1B-HF
Wikitext PPL 1.7151 1.7016 1.7142 1.5807 1.5798
Tput (4k/1k) 9306 6218 11283 9306 /
Tput (2k/1k) 11948 8961 13890 11948 7486
Model Architectural Config
$n_{\text{layers}}$ 16 16 16 16 16
$d_{\text{model}}$ 2048 2560 2560 2048 2048
$r_{\text{mlp}/\text{attn}}$ 4.8 1.067 3.6 4.8 3
GQA 4 4 9 4 1
$N_{\text{non-embed}}$ 973M 975M 965M 973M 1.074B
Table 8: Comparison against open-source models at the 3B scale. We compare our pretrained LLaMA-3.2-3B, Panda-3B, and Surefire-3B models with LLaMA-3.2-3B-HF and Qwen2.5-3B-HF in terms of inference throughput (on H200 GPUs using vLLM) and byte-level WikiText perplexity.
Model LLaMA-3.2-3B Panda-3B Surefire-3B LLaMA-3.2-3B-HF Qwen2.5-3B-HF
Wikitext PPL 1.6489 1.6454 1.6462 1.5164 1.6185
Tput (4k/1k) 4311 3335 4842 4311 6470
Model Architectural Config
$n_{\text{layers}}$ 28 28 28 28 36
$d_{\text{model}}$ 3072 4096 4096 3072 2048
$r_{\text{mlp}/\text{attn}}$ 3 1 1 3 7.167
GQA 3 3 7 3 8
$N_{\text{non-embed}}$ 2.82B 2.82B 2.82B 2.82B 2.77B
Our observations are as follows:
  • OLMo-2-1B-HF is relatively close to our predicted optimal design, with an MLP-to-attention ratio of 3 (near our predicted 3.6), but remains inference-inefficient due to its hidden dimension and GQA choices.
  • At the 3B scale, LLaMA-3.2-3B-HF achieves good accuracy but is not inference-efficient, while Qwen2.5-3B-HF is inference-efficient but less accurate.
These comparisons further underscore the necessity and relevance of our inference-efficient, high-accuracy model designs.
Appendix I Additional Results: Loss vs. GQA
We analyze the relationship between training loss and GQA while fixing the number of parameters, hidden size, and MLP-to-Attention ratio. In order to keep the number of attention parameters fixed, we vary GQA by holding the total number of heads fixed (i.e., total heads = query-heads + KV-heads) and re-allocating this fixed budget between query and key–value heads, producing asymmetric changes in their effective dimensionalities.
As shown in Figure 24, unlike hidden size and MLP-to-Attention ratio, the relationship between loss and GQA is highly fluctuating. Varying GQA does not adjust model capacity in the coordinated way that changing $d_{\text{model}}$ or $r_{\text{mlp/attn}}$ does, where the dimensions of query, key, and value scale together predictably. Specifically, note the following facts when the total number of heads is fixed
  • Increasing the number of query-heads expands the query projection dimensionality but simultaneously reduces the number of KV-heads, increasing KV sharing and thus reducing KV expressivity.
  • Conversely, decreasing query-heads increases KV-head capacity (fewer replicas) but reduces the projection dimensionality of both query and KV.
These opposing effects create a tradeoff, making the relationship between GQA and training loss non-smooth and often highly fluctuating.
Prior work shows only that query and KV projections can have non-interchangeable roles (e.g., head-importance heterogeneity Voita et al. (2019)), but provides no monotonic or predictive theory for how reallocating capacity across query versus KV should affect loss. Consistent with this, recent open LLMs choose different GQA settings even within a single family: Qwen3 uses GQA = 2 for 0.6B/1.7B, GQA = 4 for 4B/8B, GQA = 5 for 14B, GQA = 8 for 32B and for the 30B-A3B MoE; LLaMA-3/3.1/3.2 likewise use GQA = 4, 8, and 3 across closely related sizes. This variation across models of similar architecture shows that GQA is treated as a discrete, model-specific hyperparameter, supporting our decision to tune it via local search rather than integrate it into the continuous scaling law.